solve for the variable:
3(a+4)+a = 7a-(a-2)-10

Answer by TimothyLamb(4379)

By algebraic simplification:
---
3(a + 4) + a = 7a - (a - 2) - 10
3a + 12 + a = 7a - a + 2 - 10
4a + 12 = 6a - 8
6a - 4a = 12 + 8
2a = 20
---
Answer:
a = 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php